OpenStudy (anonymous):

solve and check. t^2+12t+20=0 t=

OpenStudy (anonymous):

First you factor the equation

hero (hero):

t^2+10t + 2t + 20 = 0 t(t+10)+2(t+10) =0 (t+10)(t+2) = 0 t = -2,-10

OpenStudy (anonymous):

Thanks hero!